OpenStudy (anonymous):

one side of a rectangle is 20cm. if the other side is decreased by 6cm the diagonal will be decreased by 4cm. find the length of the other side.

hartnn (hartnn):

let other side of the rectangle be x. if other side decreased by 6 ---> x-6 u know how to find the diagonal of rectangle ??

OpenStudy (anonymous):

[x^{2} +20^{2}=y^{2}] y=original length of the diagonal (x-6)^2+20^2=(y-4)^2

hartnn (hartnn):

so lets isolate one variable from those 2 equations....

hartnn (hartnn):

y from 1st equation is\(\sqrt{x^2+400}\) substitute in 2nd eq.: \((x-6)^2+400=(\sqrt{x^2+400}-4)^2\) evaluate Right side first and tell what u get...?

hartnn (hartnn):

okk,u get \(x^2-12x+36+400=x^2+400+16-8\sqrt{x^2+400}\) to get \(2\sqrt{x^2+400}=3x-5\) now squaring again: \(4(x^2+400)=9x^2+25-30x\) solving this u get,x=21. ok?

hartnn (hartnn):

the other negative root x=-15 is ignored because the length cannot be negative.

OpenStudy (anonymous):

yap.. thanks much:)

hartnn (hartnn):

welcome :)